A 0.69 −mm -diameter copper wire carries a tiny current of 2.0 μA .
Answer #1

drift speed is v = i/(n*e*A)


i = 2uA

n is the no.of free electrons per unit volume = 8.47*10^28 /m^3

e is the charge of the electron = 1.6*10^-19 C

A is the area of cross section pi*(d/2)^2 = 3.142*(0.69/2)^2*10^-6
= 0.3739*10^-6 m^2
then vd = (2*10^-6)/(8.47*10^28*1.6*10^-19*0.3739*10^-6) = 0.3947*10^-11 m/s = 3.947*10^-12 m/s